OpenStudy (anonymous):

How many solutions are there to: x^2-k^2=0?

OpenStudy (doc.brown):

With more unknowns than equations you have infinitely many solutions. In this case if x is 1 than k is 1 if x =2, k=2 x=n, k=n n can mean any real number

OpenStudy (anonymous):

but why would x and k have the same value if they are different variables?

OpenStudy (doc.brown):

Good question. If x = 2, what does k have to equal for us to minus it to get 0?

OpenStudy (anonymous):

oh! then if you say x is 2 and, let's say, k=3 and when you minus them, you get -1=0, right? So it has to be 0=0 for it to be infinate answers?

OpenStudy (anonymous):

*-1^2=0 i mean

OpenStudy (doc.brown):

Exactly. Or you could do some algebra and add k^2 to each side, if that helps.\[x^2-k^2=0\]\[x^2-k^2+k^2=0+k^2\]\[x^2=k^2\]

OpenStudy (anonymous):

Oh I see. Well thank you so much!

OpenStudy (anonymous):

Do you think you can answer another question I have?

OpenStudy (doc.brown):

Sure. Pick a best response, close this question, and post a new question. I will help with it.

OpenStudy (anonymous):

Ok I'll tag u :D